OpenStudy (anonymous):

Which ratio can form a proportion with 2/3

OpenStudy (texaschic101):

proportions are nothing but equivalent fractions. 2/3 = 4/6 = 6/9...etc 2/3 = 2:3 or 4:6 or 6:9...etc

OpenStudy (anonymous):

thanks